You love the donut store coffee and spend $2.25 per day to get a cup. After taking your finance course you decide that you will save that money and put it away towards retirement by depositing the coffee money to an account that pays 6 percent annual rate. Assuming you are 20 years old now and will start the deposit tomorrow, how much will you have in your bank account when you are 66 years old?
Answer #1

Formula: The Future Value of an ordinary annuity (FV)

FV= C× {[(1+r)^n]-1}/r

FV = Future value (The cummulative amount available in Future)
C= Periodic cash out flow.
r =effective interest rate for the period.
n = number of periods.

1. It is assumed that he will deposit everday for 365 days a year starting from tomorrow.

FV= 2.25× {[(1+0.0001596)^16790]-1}/0.0001596

FV = $191,530.91

FV = Future value (The cummulative amount available in Future)
C= Periodic cash out flow = 2.25
r =effective interest rate for the period. = [(1.06)^(1/365)]- 1 = 0.0001596
n = number of periods = (66-20)x365 =16,790

Retirement Money at 66 years= $191,530.91